pH calculation of mixture 16.42

a] 160.0 mL of 0.25M HF with 225.0 mL of 0.32M NaF
b] 165.0 mL of 0.11M C2H5NH2 with 285.0 mL of 0.20M C2H5NH3Cl


THANKS!
Answer #1
a) millimoles of HF = 160 x 0.25 = 40
millimoles of NaF = 225 x .32 = 72
final volume = 160 + 225 ml = 385 ml
[HF] = 40/385 M = 0.104M
[NaF] = 72/385 M = 0.187M
pKa of HF = 3.14
pH = pKa + log[NaF]/[HF]
pH = 3.14 + log(0.187/0.104)
pH = 3.14 + 0.255
pH = 3.359
answered by: Tiny